OpenStudy (anonymous):

∫ square root of 1+3x^3 (x^2)dx

OpenStudy (anonymous):

u- sub for this one. put \[u=1+3x^3, du=9x^2dx,\frac{1}{9}du=x^2\] and then integrate \[\frac{1}{9}\int\sqrt{u}du\] using the power rule backwards

OpenStudy (anonymous):

so is that the final answer?

OpenStudy (anonymous):

oh no you still have to find the "anti - derivative"

OpenStudy (anonymous):

totally lost

OpenStudy (anonymous):

you are looking for the function whose derivative is \[\sqrt{1+3x^3}x^2\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

you are going to run the chain rule backwards, because this looks almost like a composite function times its derivative

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

so you have to figure out how to adjust so that the constants work out correctly. that is the point of the u - sub

OpenStudy (anonymous):

do you know how to find the "anti - derivative" of \[\sqrt{u}\]?

OpenStudy (anonymous):

u^1/2

OpenStudy (anonymous):

that is another name for \[\sqrt{u}\] it is \[\sqrt{u}=u^{\frac{1}{2}}\] you still need the anti derivative

OpenStudy (anonymous):

you have to add one to the exponent and then divide by that number

OpenStudy (anonymous):

3/2u^3/2

OpenStudy (anonymous):

right. exactly!

OpenStudy (anonymous):

I shocked myself

OpenStudy (anonymous):

so now we have changed the original integral to \[\frac{1}{9}\int\sqrt{u}du\] and the anti derivative will be \[\frac{1}{9}\times \frac{2}{3}u^{\frac{3}{2}}\]

OpenStudy (anonymous):

2/27

OpenStudy (anonymous):

oh actually you made a slight mistake. dividing by 3/2 is the same as multiplying by 2/3, so it should have been \[\frac{2}{3}u^{\frac{3}{2}}\]

OpenStudy (anonymous):

oh yeah forgot to flip

OpenStudy (anonymous):

in any case you get \[\frac{2}{27}u^{\frac{3}{2}}\] and now replace u by \[1+3x^3\] to get \[\frac{2}{27}(1+3x^3)^{\frac{3}{2}}\]

OpenStudy (anonymous):

or if you prefer \[\frac{2}{27}\sqrt{1+3x^3}^3+C\]

OpenStudy (anonymous):

what about the x^2

OpenStudy (anonymous):

you can check your answer by taking the derivative and seeing that you will get \[\sqrt{1+3x^3}\times x^2\]

OpenStudy (anonymous):

the chain rule will give it to you . that was the idea behind the original substitution, making \[u=1+3x^3\]

OpenStudy (anonymous):

ummmm???

OpenStudy (anonymous):

let us do a simpler example. what is the derivative of \[\sin(x^2)\]?

OpenStudy (anonymous):

OH I get it that's wehre the 1/9 came from

OpenStudy (anonymous):

righto!~

OpenStudy (anonymous):

duh...okay thank you fro helpng me to the answer 2/27 SQRT(1+3x^3 +C

OpenStudy (anonymous):

careful it is \[\frac{2}{27}\sqrt{1+3x^3}^3+C\]

OpenStudy (anonymous):

don't forget the "cube" because you are raising to the power of 3/2

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

Can you walk me through another one? lim e^3x-1/4x x->0